OpenStudy (anonymous):

One half of the distance between the nuclei of two adjacent metal atoms in a close crystal lattice is called

OpenStudy (asevilla5):

Two adjacent metal atoms' nuclei have a distance which is simply called the 'metallic atomic radius'. This named radius is different from that of non-metallic or gaseous atomic structures.

OpenStudy (anonymous):

@asevilla, thanks

OpenStudy (anonymous):

i need a help, solve the set of linear equations by the matrix method : a+3b+2c=3 , 2a-b-3c= -8, 5a+2b+c=9. Sove for c

OpenStudy (anonymous):

Solve the set of linear equations by Guassian elimination method : a+2b+3c=5, 3a-b+2c=8, 4a-6b-4c=-2. Find c